Carnes Cosmetics Co.'s stock price is $68.74, and it recently paid a $2.50 dividend. This dividend is expected to grow by 28% for the next 3 years, then grow forever at a constant rate, g; and rs = 16%. At what constant rate is the stock expected to grow after Year 3? Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

Value of stock is equal to the present value of all future dividends

= Expected dividend/(required return - growth rate)

68.74 = 2.50(1.28)/(1.16) + 2.50(1.28)^2/(1.16)^2 + 2.50(1.28)^3/(1.16)^3 + 2.50(1.28)^3(1+g)/(16%-g)(1.16)^3

68.74 = 2.7586+3.0440+ 3.3589 + 3.3589(1+g)/(16%-g)

59.5785 = 3.3589(1+g)/(16%-g)

17.7375 = (1+g)/(0.16-g)

2.838-17.7375g = 1+g

18.7375g = 1.838

G = 9.81%

Hence, constant growth rate rate after year 3 = 9.81%